The student body at St. Algebra College attends an average of 3.3 parties per month. A random sample of 117 sociology majors averages 3.8 parties per month, with a standard deviation of 0.53. Are sociology majors significantly different from the student body as a whole?
1
Expert's answer
2012-10-26T11:21:44-0400
Build the 95% confidence interval for sociology major’s average number of parties.
Z(1 – (1 – 0.95)/2) = z(0.975) = 1.96
Margin of error:
ME = z * SD / sqrt(n) = 1.96*0.53/sqrt(117) ≈ 0.096
The confidence interval looks:
CI = (P – ME, P + ME) = (3.8 – 0.096, 3.8 + 0.096) ≈ (3.7, 3.9)
The average for all St. Algebra College students, 3.3, is not inside the interval (3.7, 3.9), so under 95% confidence level sociology majors are significantly different from the student body as a whole.
For the 99% confidence level we will get the same result.